OpenStudy (anonymous):

Which of the following is one of the factors of x2 – 13x + 12?

OpenStudy (anonymous):

is there a list to choose from? ^_^

OpenStudy (anonymous):

(x-1)(x-12)

OpenStudy (anonymous):

this factors to (x - 12)(x - 1)

OpenStudy (anonymous):

x^2 -13x+12=0 (x-1)(x-12)=0 so x= 1, 12

OpenStudy (anonymous):

x^2-12x-x+12 = x(x-12)-1(x-12) =(x-1)(x-12)

OpenStudy (anonymous):

.-. I think you've got your answer now dav ^_^

OpenStudy (anonymous):

x + 1 x – 1 x + 12 x – 3

OpenStudy (anonymous):

x-1